Circular No. 7625

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 J1 (NEAT) MPEC 2001-J34 contains 40 astrometric observations May 11–15 and

parabolic orbital elements (T = 2001 Mar. 19, q = 1.00 AU, ω = 279^{o}, Ω = 198^{o}, i = 11^{o}, equinox 2000.0) computed from 26 of them. The orbit is still very indeterminate, and it is not clear whether or not the comet is of short period.

COMET C/2001 A2 (LINEAR) J. Broughton, Reedy Creek, Queensland, writes that his CCD obser-

vations (0.25-m f/6.6 Schmidt-Cassegrain) on May 9.3 UT showed the brighter (western) component at m_{2} = 14.7, with the secondary, 0.3 mag fainter, perhaps 5′′ distant in p.a. 100^{o}, although some elongation of the secondary could make the separation as much as 7′′. On May 10.3 the appearance was similar. On May 11.3 the western component had brightened to m_{2} = 13.2, and the eastern component had virtually disappeared. On May 12.4 the western component had faded to m_{2} ∼ 14.0, while the secondary was then perhaps 2 mag fainter, 8′′.5 away in p.a. 100^{o}.

Using the absolute astrometry for Apr. 30 (see IAUC 7616) and May 9 given on MPC 42656 (where the western primary is labeled B and the eastern secondary is labeled A), Z. Sekanina, Jet Propulsion Laboratory, obtained a separation time of Mar. 17 ± 12 and an acceleration of 7.1 ± 2.4 (in units of 10^{−5} solar gravity).

Total visual magnitude estimates by A. Pearce, Nedlands, Western Australia (20 × 80 binoculars): May 9.48 UT, 6.2; 11.44, 5.3; 13.49, 5.2; 14.49, 5.4; 15.49, 5.4.

SUPERNOVA 2001bf IN MCG +04-42-22 H. Kawakita and K. Kinugasa, Gunma Astronomical Observatory; and

H. Yamaoka, Kyushu University, write: “Despite the low-S/N ratio, a lowresolution spectrum of SN 2001bf taken with the Gunma 0.65-m telescope (+ GCS spectrograph) on May 11.61 UT clearly shows a deep and broad Si II absorption feature around 610 nm on a rather blue continuum, indicating that this is a type Ia supernova shortly before maximum light. Other broad absorption features typical of type Ia events are also seen.”

Visual magnitude estimates by J. D. West, Mulvane, KS: May 7.44 UT, 14.6; 8.44, 14.4. An unfiltered CCD measurement, also by West, on May 14.438 gave R = 14.5 ± 0.2.

2001 May 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden